OpenStudy (just_one_last_goodbye):

Choose the equation below that represents the line passing through the point (-2, -3) with a slope of -6.

OpenStudy (just_one_last_goodbye):

A. y + 3 = -6(x + 2)

OpenStudy (just_one_last_goodbye):

B.y + 2 = 6(x+3)

OpenStudy (just_one_last_goodbye):

C. y - 3 = -6(x-2)

OpenStudy (just_one_last_goodbye):

D. y - 2 = 6(x-3)

OpenStudy (anonymous):

Use the point slope formula

OpenStudy (anonymous):

y-y1=m(x-x1)

OpenStudy (just_one_last_goodbye):

Will you guide me through the steps?

OpenStudy (anonymous):

m=slope

OpenStudy (anonymous):

so what will you equation look like when the slope is inserted?

OpenStudy (anonymous):

y-y1=slope(x-x1) plug it in

OpenStudy (just_one_last_goodbye):

Think its C or D

OpenStudy (anonymous):

show me the work.

OpenStudy (anonymous):

Do not guess.

OpenStudy (just_one_last_goodbye):

your not helping im new at this

OpenStudy (anonymous):

(x1,y1) m=slope that is all you need to know

OpenStudy (anonymous):

Then just plug it all in \[y-y1=m(x-x1) \]

OpenStudy (just_one_last_goodbye):

Any hints?

OpenStudy (anonymous):

Well everything you need to know is present you now just need to plug it in there is nothing else to this.

OpenStudy (just_one_last_goodbye):

Ok lets see give me a second to start thinking.

OpenStudy (just_one_last_goodbye):

Question which is the slope? is it -6 or the coordinates?

OpenStudy (anonymous):

m=slope Choose the equation below that represents the line passing through the point (-2, -3) with a slope of -6.

OpenStudy (just_one_last_goodbye):

Ah ok -6.

OpenStudy (anonymous):

Use this website to help you learn https://www.mathsisfun.com/algebra/line-equation-point-slope.html

jabez177 (jabez177):

I think that guy has a good point. I hate these types. I can learn from that website as well. See ya!

jabez177 (jabez177):

Time to get to work.